2 angles are complimentary. The measure of the larger angle is 4 more than 3 times the smaller angle. Write an equation and solve to find the measure of the angles.
Answer by ReadingBoosters(3246)

complementary angles have a sum of 90
x + y = 90
x = 4 + 3y
Substitute into equation
(4 + 3y) + y = 90
4y = 86
y = 21.5 degrees
x = 4 + 3(21.5) = 68.5 degrees